ROTATIONAL ANALYSIS OF THE C ^{2}Π STATE OF SrF BY OPTICAL-OPTICAL DOUBLE RESONANCE SPECTROSCOPY

\underline{P. M. SHERIDAN}, J.-G. WANG, Department of Chemistry, University of Waterloo, 200 University Ave. West, Waterloo, ON, N2L 3G1 Canada; M. J. DICK, Department of Physics, University of Waterloo, 200 University Ave. West, Waterloo, ON, N2L 3G1 Canada; and P. F. BERNATH, Department of Chemistry, University of Waterloo, 200 University Ave. West, Waterloo, ON, N2L 3G1 Canada.

The C ^{2}Π (v = 0) state of SrF has been investigated at high resolution using the technique of optical-optical double resonance spectroscopy. SrF was synthesized in a Broida-type oven by the reaction of Sr metal vapor with SF_{6}. The band heads of the A ^{2}Π−X ^{2}Σ^{+} transition were first excited using a linear dye laser (∼1 cm^{−1} bandwidth). A Ti:Sapphire ring laser was then utilized to record high-resolution laser excitation spectra of the C ^{2}Π_{1/2} − A ^{2}Π_{1/2} and C ^{2}Π_{3/2} − A ^{2}Π_{3/2} transitions. A rotational analysis of these transitions is in progress. In addition, an investigation of the D ^{2}Σ^{+} state is currently underway and preliminary observations will be presented.